Wersja testu A 14 lutego 2005r.

1. a. N b. N c. T d. N 2. a. N b. T c. N d. T 3. a. T b. N c. T d. N 4. a. T b. T c. T d. N 5. a. T b. T c. N d. N 6. a. N b. N c. N d. T 7. a. N b. T c. T d. T 8. a. T b. T c. N d. N 9. a. N b. T c. N d. T 10. a. T b. T c. T d. N 11. a. T b. T c. T d. N 12. a. N b. N c. T d. T 13. a. T b. T c. N d. T 14. a. T b. N c. N d. T 15. a. N b. N c. T d. T 16. a. N b. T c. N d. T 17. a. T b. N c. N d. N 18. a. T b. T c. N d. N 19. a. T b. N c. T d. T 20. a. N b. N c. T d. T

Wersja testu B 14 lutego 2005r.

1. a. N b. N c. T d. N 2. a. N b. T c. T d. N 3. a. T b. T c. N d. N 4. a. T b. T c. N d. T 5. a. T b. T c. N d. N 6. a. N b. N c. N d. T 7. a. T b. T c. T d. N 8. a. T b. N c. N d. T 9. a. T b. N c. N d. T 10. a. T b. T c. T d. N 11. a. T b. N c. T d. T 12. a. T b. N c. T d. N 13. a. T b. N c. T d. T 14. a. N b. T c. N d. T 15. a. T b. N c. T d. N 16. a. T b. T c. N d. N 17. a. N b. N c. T d. N 18. a. N b. T c. T d. N 19. a. T b. T c. N d. T 20. a. N b. N c. T d. T

Wersja testu C 14 lutego 2005r.

1. a. T b. N c. N d. N 2. a. N b. N c. T d. T 3. a. T b. N c. N d. T 4. a. T b. N c. T d. T 5. a. N b. T c. T d. N 6. a. N b. T c. N d. N 7. a. T b. T c. N d. T 8. a. T b. T c. N d. N 9. a. N b. T c. T d. N 10. a. T b. T c. N d. T 11. a. N b. T c. T d. T 12. a. N b. T c. T d. N 13. a. T b. T c. N d. T 14. a. N b. T c. N d. T 15. a. T b. N c. N d. T 16. a. N b. T c. N d. T 17. a. N b. N c. N d. T 18. a. N b. T c. T d. N 19. a. N b. T c. T d. T 20. a. T b. N c. N d. T

Wersja testu D 14 lutego 2005r.

1. a. N b. N c. T d. N 2. a. T b. N c. T d. N 3. a. N b. T c. N d. T 4. a. T b. T c. N d. T 5. a. N b. N c. T d. T 6. a. N b. N c. N d. T 7. a. T b. T c. N d. T 8. a. T b. T c. N d. N 9. a. N b. T c. T d. N 10. a. T b. N c. T d. T 11. a. T b. N c. T d. T 12. a. N b. N c. T d. T 13. a. N b. T c. T d. T 14. a. T b. N c. N d. T 15. a. T b. N c. N d. T 16. a. N b. T c. T d. N 17. a. N b. N c. T d. N 18. a. T b. N c. T d. N 19. a. T b. T c. N d. T 20. a. T b. N c. T d. N